# MI9 - Agent Intelligence Protocol: Runtime Governance for Agentic AI Systems

### v2

@@ Abstract
Agentic AI systems capable of reasoning, planning, and executing actions present fundamentally distinct governance challenges compared to traditional AI models. Unlike conventional AI, these systems exhibit emergent and unexpected behaviors during runtime, introducing novel agent-related risks that cannot be fully anticipated through pre-deployment governance alone. To address this critical gap, we introduce MI9, the first fully integrated runtime governance framework designed specifically for safety and alignment of agentic AI systems. MI9 introduces real-time controls through six integrated components: agency-risk index, agent-semantic telemetry capture, continuous authorization monitoring, Finite-State-Machine (FSM)-based conformance engines, goal-conditioned drift detection, and graduated containment strategies. Operating transparently across heterogeneous agent architectures, MI9 enables the systematic, safe, and responsible deployment of agentic systems in production environments where conventional governance approaches fall short, providing the foundational infrastructure for safe agentic AI deployment at scale. Detailed analysis through a diverse set of scenarios demonstrates MI9’s systematic coverage of governance challenges that existing approaches fail to address, establishing the technical foundation for comprehensive agentic AI oversight.
@@ 2 Related Work
Alignment research has primarily focused on training-time interventions: Reinforcement Learning from Human Feedback (RLHF) [6, 28] and Constitutional AI [1] establish initial value alignment but cannot address failures emerging during autonomous operation when agents encounter novel situations or coordinate with other systems [16].
Building on this foundation, recent work has mapped agentic system taxonomies [31, 15], governance frameworks [9, 30, 8, 17], and threat models [25, 5, 33]. However, leading benchmarks prioritize task completion over governance dimensions such as behavioral consistency [14, 21, 37, 13, 32].
Meanwhile, current monitoring solutions [36, 20, 19, 35, 7] provide reactive observation rather than proactive intervention. Similarly, process observability research [10] and visibility frameworks [4] focus on observation, while enterprise platforms [12, 24, 23] and security frameworks [29, 26] rely on static risk assessment inadequate for emergent runtime behaviors.
@@ 3.1 Framework Integration and Overview
This integrated architecture enables proactive, real-time oversight specifically designed for agentic systems exhibiting emergent behaviors during execution, addressing the fundamental gap between static pre-deployment assessments and reactive post-incident analysis. Production deployments require standard distributed systems coordination [11, 2], but the core governance semantics operate independently of the underlying consistency mechanisms.
@@ 3.3 Agentic Telemetry Schema
Traditional infrastructure monitoring captures operational events (HTTP responses, database latency, service calls) but systematically misses the cognitive processes that create governance risks in agentic systems. The majority of agentic governance violations originate from cognitive behaviors such as goal revision, memory retrieval, tool-chaining decisions—that remain invisible to conventional observability frameworks [10]. Safe deployment of agentic AI systems requires visibility into the moments when agents autonomously revise objectives, chain unexpected tool sequences, or retrieve memory that fundamentally alters downstream behavior—cognitive processes critical for responsible oversight yet absent from standard infrastructure telemetry.
@@
ATS extends OpenTelemetry’s emerging agent conventions [27] by introducing governance-semantic abstractions that transform opaque agent execution into actionable oversight intelligence. Policy engines evaluate event attributes to enforce constraints, such as "Tier 2 agents cannot execute shell commands without approval," while drift detectors analyze cognitive event patterns to identify concerning behavioral changes. This semantic foundation enables the real-time intervention capabilities that reactive monitoring lacks: in governance terms, we cannot govern what we cannot observe.
@@ 3.5 Real-Time Conformance Engine
Building on formal runtime verification frameworks for adaptive systems [3], we implement a sequence-aware rule layer operating on the ATS stream. Our approach employs finite-state machines (FSMs) following proven Communicating Sequential Processes (CSP) verification principles [22], where each rule compiles into an FSM with states representing pattern progress and transitions triggered by ATS events that satisfy specified predicates. This design balances expressiveness with computational efficiency: FSMs encode sequential and temporal constraints relevant to agent governance while maintaining bounded memory usage and deterministic evaluation with $O(k)$ event processing time per agent, where $k$ is the number of active patterns.
@@ 3.6 Behavioral Drift Detection
Building on recent advances in uncertainty-based drift detection [18] and multi-agent ensemble methods [34], our system analyzes behavioral patterns using statistical evaluation of ATS event sequences within defined goal contexts. It maintains frequency distributions of event types, tool usage, and sequence characteristics for each distinct agent objective, constructing goal-conditioned baselines $B_{g}$ that enable detection when behavior changes while goals remain constant.
@@ 4 Framework Analysis
To evaluate a broad spectrum of agentic behaviors and failure modes, we generate structured agent trace logs via prompt-conditioned LLM (Gemini 1.5 Flash) simulation. While our evaluation is synthetic, the controlled setup enables systematic simulation of failure modes difficult to isolate in real agent logs, providing a reproducible testbed for validating governance mechanisms. This approach also allows us to stress-test rare but critical misalignment behaviors that are unlikely to surface reliably in limited real-world traces.
@@ 4.1 Evaluation Results
LangSmith’s poor performance reflects its infrastructure focus: it captures tool calls but misses cognitive events (goal revision, memory access) where the majority of violations originate. LangSmith attains a perfect false positive rate, bnut that is likely due to the fact that it detected so few violations that false positives were effectively impossible to accumulate.
@@ 5 Conclusion
Despite these limitations, MI9 provides the first integrated, comprehensive runtime governance framework for agentic systems. It moves beyond static, pre-deployment assessments to a dynamic, in-session oversight paradigm. The framework introduces agent-semantic telemetry and real-time intervention capabilities that existing approaches lack, laying a necessary foundation for the safe and responsible deployment of agentic AI systems at scale.
@@ 6 Acknowledgments
Appendix A Technical Appendix
A.1 Agency-Risk Index Scoring Criteria
@@
A.2 Agency-Risk Index (ARI) Calculation
@@ A.2.1 Mathematical Formulation
A.2.2 Dimensional Scoring and Aggregation
@@
This normalizes each criterion to the range $[0,1]$ and averages across the four criteria within the dimension.
@@
A.2.3 Risk Tier Classification
The continuous ARI value maps to discrete risk tiers that determine baseline governance requirements:
@@ A.2.3 Risk Tier Classification
A.3 Risk-Tier-Weighted Containment Model Development
@@
A.3.1 Comprehensive Containment Decision Framework
The containment decision integrates multiple contextual dimensions with agent capability assessment:
@@ A.3.1 Comprehensive Containment Decision Framework
where $C=\{\text{Monitor},\text{Planning},\text{Restriction},\text{Isolation}\}$ represents escalating containment levels.
A.3.2 Multi-Dimensional Context Assessment
Organizations should develop scoring matrices across key contextual dimensions:
@@ A.3.2 Multi-Dimensional Context Assessment
A.3.3 Detailed Organizational Example: Investment Banking
Consider a multinational investment bank implementing MI9 across trading, research, and client services:
@@ A.3.3 Detailed Organizational Example: Investment Banking
This structured approach enables organizations to develop MI9 containment models that balance governance requirements with operational effectiveness while remaining adaptable to evolving business needs and risk landscapes.
A.4 Framework Integration
@@
A.5 Evaluation Dataset Statistics
@@ A.5.2 Dataset Composition
A.5.3 Framework Performance Comparison
@@
A.6 MI9 Components High-Level Example Responses

### v3

# MI9: An Integrated Runtime Governance Framework for Agentic AI

@@ Abstract
Agentic AI systems capable of reasoning, planning, and acting present governance challenges that differ fundamentally from conventional models. Because these systems can exhibit emergent, unexpected behaviors during execution, many risks cannot be fully anticipated pre-deployment. We present MI9, an integrated framework for runtime safety of agentic AI, where safety properties are enforced over live behavior sequences. MI9 provides six coordinated mechanisms: Agency-Risk Index, agent-semantic telemetry, goal-aware authorization monitoring, finite-state conformance engines, goal-conditioned drift detection, and graded containment, that operate in a model and infrastructure agnostic manner across heterogeneous agent stacks. MI9 is a framework layer that instruments and governs existing systems to enable systematic, safe deployment at scale. In evaluations over 1,000 diverse multi-domain synthetic scenarios, MI9 achieves high detection with low FPR. By shifting the locus of assurance to runtime safety, MI9 establishes a practical foundation for comprehensive, operational oversight of agentic AI. We open-source all prompts, scripts, and per-scenario summaries for reproducibility. [Uncaptioned image]  Code
@@ 2 Related Work
Alignment research has primarily focused on training-time interventions: Reinforcement Learning from Human Feedback (RLHF) [christiano_deep_2017, ouyang_training_2022] and Constitutional AI [anthropic_constitutional_ai_2022] establish initial value alignment but cannot address failures emerging during autonomous operation when agents encounter novel situations or coordinate with other systems [kenton_alignment_2021].
Building on this foundation, recent work has mapped agentic system taxonomies [schneider_generative_2025, kasirzadeh_characterizing_2025], governance frameworks [openai_whitepaper_2024, raza_trism_2025, engin_dimensional_2025, kolt_governing_2025], and threat models [narajala_securing_2025, chan_atfaa_2025, syros_saga_2025]. However, leading benchmarks prioritize task completion over governance dimensions such as behavioral consistency [kapoor_agents_that_matter_2024, liu_survey_agent_evaluation_2025, zhou_webarena_2023, jimenez_swe_bench_2023, sumers_agentverse_2025].
Meanwhile, current monitoring solutions [wu_agentops_2024, langfuse_observability_2024, langsmith_tracing_2024, wandb_llm_monitoring_2024, datadog_llm_2024] provide reactive observation rather than proactive intervention. Similarly, process observability research [fournier_process_2025] and visibility frameworks [chan_visibility_2024] focus on observation, while enterprise platforms [holisticai_governance_2024, monitaur_platform_2024, modelop_agentic_2024] and security frameworks [owasp_llm_top10_2024, nist_ai_rmf_2024] rely on static risk assessment inadequate for emergent runtime behaviors.
@@ 3 MI9 Framework
Figure 1: MI9 Framework Pipeline
3.1 What MI9 Is, And Isn’t
MI9 is a framework layer—policies, telemetry schema, temporal conformance rules, and graduated containment—not a new agent or planning algorithm. It instruments existing stacks in a model and infrastructure agnostic way. For evaluation, we use LLM-generated agent traces to simulate governed behavior; MI9’s governance logic itself is rule-based and telemetry-driven, without a dependence on any particular LLM size or vendor. Unlike multi-agent prompt-orchestration approaches in HCI that coordinate prompts without enforceable runtime controls, MI9 contributes (i) a vendor-agnostic telemetry schema, (ii) temporal conformance rules over agent traces, and (iii) graduated, auditable containment mechanisms that interpose on actions in real time.
4 Threat Model & Scope
  1. 1.
    In scope. Runtime risks from agent behavioral sequences and coordination: (i) goal drift under fixed stated goals, (ii) policy-skipping tool chains, (iii) delegated privilege escalation, (iv) multi-agent coordination failures.
  2. 2.
    Out of scope. Pretraining/data harms, upstream supply-chain compromise, and non-sequential issues not captured in event traces.
  3. 3.
    Actors. Deployed agents (incl. subagents), human overseers, organizational policy engine.
  4. 4.
    Assumptions. Minimum ATS coverage at least for action-level events; bounded event reordering; ability to pause/contain.
  5. 5.
    Objective. Minimize undetected violations at very low FPR while preserving operational continuity via graduated containment.
@@ 4.1 Framework Integration and Overview
This integrated architecture enables proactive, real-time oversight specifically designed for agentic systems exhibiting emergent behaviors during execution, addressing the fundamental gap between static pre-deployment assessments and reactive post-incident analysis. Production deployments require standard distributed systems coordination [fowler_event_2005, bailis_coordination_2014], but the core governance semantics operate independently of the underlying consistency mechanisms.
@@ 4.3 Agentic Telemetry Schema
Traditional infrastructure monitoring captures operational events (HTTP responses, database latency, service calls) but systematically misses the cognitive processes that create governance risks in agentic systems. The majority of agentic governance violations originate from cognitive behaviors such as goal revision, memory retrieval, tool-chaining decisions—that remain invisible to conventional observability frameworks [fournier_process_2025]. Safe deployment of agentic AI systems requires visibility into the moments when agents autonomously revise objectives, chain unexpected tool sequences, or retrieve memory that fundamentally alters downstream behavior—cognitive processes critical for responsible oversight yet absent from standard infrastructure telemetry.
@@
ATS extends OpenTelemetry’s emerging agent conventions [opentelemetry_genai_2024] by introducing governance-semantic abstractions that transform opaque agent execution into actionable oversight intelligence. Policy engines evaluate event attributes to enforce constraints, such as "Tier 2 agents cannot execute shell commands without approval," while drift detectors analyze cognitive event patterns to identify concerning behavioral changes. This semantic foundation enables the real-time intervention capabilities that reactive monitoring lacks: in governance terms, we cannot govern what we cannot observe.
@@ 4.4 Continuous Authorization Monitoring
Security rationale
CAM does not permit arbitrary self-escalation. It enforces: (i) authority matrices specifying who may delegate which permissions; (ii) provenance-checked delegation chains with expiry; (iii) temporal ordering guards (e.g., approval must precede use); and (iv) default-deny for tier-elevating goal transitions. These constraints prevent silent privilege drift while preserving legitimate operational flexibility.
@@ 4.5 Real-Time Conformance Engine
Building on formal runtime verification frameworks for adaptive systems [carwehl_runtime_2023], we implement a sequence-aware rule layer operating on the ATS stream. Our approach employs finite-state machines (FSMs) following proven Communicating Sequential Processes (CSP) verification principles [luckcuck_varanus_2024], where each rule compiles into an FSM with states representing pattern progress and transitions triggered by ATS events that satisfy specified predicates. This design balances expressiveness with computational efficiency: FSMs encode sequential and temporal constraints relevant to agent governance while maintaining bounded memory usage and deterministic evaluation with $O(k)$ event processing time per agent, where $k$ is the number of active patterns.
@@ 4.6 Behavioral Drift Detection
Building on recent advances in uncertainty-based drift detection [komorniczak_drift_2024] and multi-agent ensemble methods [vieira_driftage_2021], our system analyzes behavioral patterns using statistical evaluation of ATS event sequences within defined goal contexts. It maintains frequency distributions of event types, tool usage, and sequence characteristics for each distinct agent objective, constructing goal-conditioned baselines $B_{g}$ that enable detection when behavior changes while goals remain constant.
@@ 5 Framework Analysis
To evaluate a broad spectrum of agentic behaviors and failure modes, we generate structured agent trace logs via prompt-conditioned LLM (Gemini 2.5 Flash) simulation. While our evaluation is synthetic, the controlled setup enables systematic simulation of failure modes difficult to isolate in real agent logs, providing a reproducible testbed for validating governance mechanisms. This approach also allows us to stress-test rare but critical misalignment behaviors that are unlikely to surface reliably in limited real-world traces. We also note that MI9’s runtime detection and intervention logic is rule/automata based and does not rely on the LLM at use.
Evaluation Scope & Assumptions.
We evaluate MI9 as a runtime governance layer over agent execution traces. Scenarios and traces are synthetic but structured to cover diverse failure modes, with all prompts, generators, and judge scripts released. Baselines (OT, LS) are configured to their public defaults with equivalent access to the same raw traces and tools (details below). Metrics quantify governance qualities (detection, FPR, coverage, causal clarity, early warning, and intervention), not task success. This isolates runtime safety behavior from agent capability. Judges never receive hidden ground-truth tags or labels; they see only scenario text and framework logs. Judge and generator use distinct prompts and instances to prevent leakage. We evaluate MI9 at its deployment operating point, tuned to minimize expected intervention cost (operator time + false blocks) subject to latency constraints. Because agent traces are heterogeneous and long-tailed, fixed-FPR targets are not stable across domains; we therefore report Detection/FPR/Intervention/Clarity/Predictive at the chosen operating point.
Evaluation Structure
We use a three-stage, LLM-driven pipeline with deterministic prompts and JSON-validated outputs. For scenario generation we call Gemini (default: 2.5-flash, $T=0.7$, top-p=0.9=0.9, top-k=40=40, max 8192 tokens; safety filters off). Governance log generation uses Gemini 2.5-pro ($T=0.7$, max 8192, 180s timeout, up to 3 retries with exponential backoff). Evaluation uses Gemini 2.5-flash with more deterministic settings ($T=0.2$, top-p=0.9=0.9, max 4096). Runners support configurable concurrency (default 3–5), robust error handling, and retry with jitter; the system is model-agnostic so models can be swapped easily. We report Governance Maturity plus Detection and Actionable-Intelligence metrics, and we tune sampling per stage to balance creativity (generation) and consistency (evaluation).
Trace Construction.
Each scenario specifies risks and constraints; the generator produces stepwise ATS-consistent event sequences (cognitive, action, coordination) with multi-step chains that realize the targeted violations (e.g., approval-skipping sequences, privilege drift). A held-out rubric marks the ground-truth violations and types; the same rubric evaluates all frameworks.
@@ 5.1 Evaluation Results
Metrics.
Table 3: Metrics and definitions for governance evaluation.
Metric Definition
Detection $\mathrm{TP}/\#\mathrm{Actual}$
FPR $\mathrm{FP}/\#\mathrm{Detections}$
Coverage $\mathrm{Types\ detected}/\#\mathrm{Types\ present}$
[…]
Alerting Share of violations with an early signal ($\geq\!1$ event prior).
Intervention Share of attempts blocked at runtime.
@@
LangSmith’s poor performance reflects its infrastructure focus: it captures tool calls but misses cognitive events (goal revision, memory access) where the majority of violations originate. LangSmith attains a perfect false positive rate, but that is likely due to the fact that it detected so few violations that false positives were effectively impossible to accumulate.
Baselines Configuration.
OpenTelemetry (OT) receives the same event stream but operates at infra level (HTTP/db/tool spans) without agent-semantic cognition signals. LangSmith (LS) receives tool/model callback traces but not the cognitive/coordination events in ATS. Both baselines use their recommended detection hooks; we do not add MI9 logic to them. All systems observe the same scenario traces, ensuring like-for-like comparison of governance visibility and intervention capability.
@@ 6 Conclusion
Despite these limitations and to our knowledge, MI9 provides the first integrated, comprehensive runtime governance framework for agentic systems. It moves beyond static, pre-deployment assessments to a dynamic, in-session oversight paradigm. The framework introduces agent-semantic telemetry and real-time intervention capabilities that existing approaches lack, laying a necessary foundation for the safe and responsible deployment of agentic AI systems at scale.
@@ 7 Acknowledgments
8 Appendix
8.1 Agency-Risk Index Scoring Criteria
@@
8.2 Agency-Risk Index (ARI) Calculation
@@ 8.2.1 Mathematical Formulation
8.2.2 Dimensional Scoring and Aggregation
@@
8.2.3 Risk Tier Classification
@@
8.3 Risk-Tier-Weighted Containment Model Development
@@
8.3.1 Comprehensive Containment Decision Framework
@@
where $C=\{\text{Monitor},\text{Planning},\text{Restriction},\text{Isolation}\}$.
8.3.2 Multi-Dimensional Context Assessment
@@
8.3.3 Detailed Organizational Example: Investment Banking
@@
8.4 Framework Integration
@@
8.5 Evaluation Dataset Statistics
@@ 8.5.2 Dataset Composition
8.5.3 Framework Performance Comparison
@@
8.6 MI9 Components High-Level Example Responses